# MultiNash-PF: A Particle Filtering Approach for Computing Multiple Local Generalized Nash Equilibria in Trajectory Games

### v2

@@ IV Identifying Local Nash Equilibria via Bayesian Inference
$\min_{\{v_{t},w_{t},\eta_{t}\}_{t=0}^{\tau}}$ $\left\lVert v_{\tau}\right\rVert_{Q_{\tau}}^{2}+\sum_{t=0}^{\tau-1}\left\lVert v_{t}\right\rVert_{Q}^{2}+\sum_{t=0}^{\tau}\left\lVert w_{t}\right\rVert_{R}^{2}+\left\lVert\eta_{t}\right\rVert^{2}_{Q_{\eta}},$
@@
where $v_{t}$, $w_{t}$, and $\eta_{t}$ are bounded disturbances, and $Q_{\eta}\in\mathbb{R}^{n_{c}\times n_{c}}$ assigns the importance of constraint satisfaction. Further, $\psi$ is a barrier function used to quantify the degree of constraint violation by defining
@@
where $\alpha$ controls the strictness enforcing constraints.
We note that the problem in (IV) resembles a moving horizon estimation problem [32]. A direct attempt at solving (IV) using traditional gradient-based methods poses a similar computational burden in identifying all of the local equilibria compared to problem (1). However, the above connection between (1) and the moving horizon estimation problem (IV) highlights the potential to treat the problem from a Bayesian inference perspective. To this end, we construct the following virtual state-space model from (IV) as
@@
with $\bar{Q}\coloneqq\mathop{\rm blkdiag}(Q^{-1},Q_{\eta}^{-1})$, $\bar{Q}_{\tau}\coloneqq\mathop{\rm blkdiag}(Q_{\tau}^{-1},Q_{\eta}^{-1})$, $\bar{R}\coloneqq\mathop{\rm blkdiag}(0_{n\times n},R^{-1})$. The state estimation relevant to (9) is to estimate the state $\bar{x}_{t}$ given the measurements $\bar{y}_{t}$. The value of the virtual measurements must be set in a way that steers the inference problem to have the same optima as in problem (1). We achieve this by noting that $y_{x,t}$ is a measurement value used to correct the state estimate of $x_{t}$ towards the reference $\hat{x}_{t}$. While $y_{g,t}$ is to satisfy constraints through (8). Observing that the barrier function in (8) outputs zero when the constraints are satisfied, we set the virtual measurement for the barrier function equal to $0_{c}$. Hence, the virtual measurement that will drive $\bar{x}_{t}$ towards optimality takes the value
@@
The Bayesian inference problem pertaining to (9) is one that characterize the conditional distribution $p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})$. In the following, lemma (1) elucidates that equivalence holds between a posteriori estimate of $p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})$ and the problem in (1).
@@
We refer the interested reader to [16, Thm. 1] for the proof. The above lemma implies that the local GNEs in (1) translate to multiple local maximums of $p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})$. Hence, the Bayesian inference problem for (9) involves characterizing the multi-modal PDF $p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})$. Generally, state estimation of a multi-modal PDF does not admit closed-form solutions [33]. Hence, we resort to approximate solutions. A powerful approach is to use Monte Carlo methods that empirically approximate the distribution via a set of $J$ particles as
@@
where $\delta(\cdot)$ is the dirac delta function and $w_{\tau}^{i}$ is the weight assigned to a sample trajectory $\bar{x}_{0:\tau}^{i}$ for $j=1,\ldots,J$.
To achieve the approximation in (10), we employ the implicit particle filter  [31] due to the twofold benefit it holds to implement the MultiNash-PF framework. First, as suggested in lemma 1, the local maximums (i.e., high-probability regions) of $p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})$ correspond to the local GNEs. Hence, focusing our sampling efforts towards these high-probability regions can reduce the number of required samples for the approximation in (10). This approach aligns precisely with the principle of implicit importance sampling, whose objective is to build a mapping hat draws particles from the high-probability regions of the filtering distribution in (10). Second, we want to avoid the traditionally needed resampling step in particle filtering methods to ensure that the inferred trajectories are smooth. Fortunately, the implicit particle filter does not suffer from the problem of particle degeneracy, and filtering can be achieved by disregarding resampling.
We leverage the Markovian property of the virtual system (9) to recursively approximate (10). Using Bayes’ rule, we have
@@
for $0\leq t\leq\tau$. This indicates that given the particle $\bar{x}_{t-1}^{j}$ from the prior distribution $p(\bar{x}_{0:t-1}\,|\,\bar{y}_{0:t-1})$, we can obtain $\bar{x}_{t}^{j}$. As shown in [31], we use implicit importamapping sampling and identify an implicit mapping to sample from high probability regions of (11). The implicit map is constructed by considering sampling from a reference random vector $\xi_{t}$ with density $p(\gamma_{t})$ and map it to high probability regions of $p(x_{0:\tau}\,|\,y_{0:\tau})$, such that a sample $\gamma^{i}_{t}$ is mapped to $x_{t}^{i}$. The mapping connects the highly probable regions of $p(\xi)$ to $p(x_{0:\tau}\,|\,y_{0:\tau})$ by letting
@@
where $F(x_{t}^{i})=-\log\left(p(x_{t}^{i}\,|\,y_{1:t})\right)$ and $G(\gamma_{t}^{i})=-\log\left(p(\gamma_{t}^{i})\right)$. Finding an explicit solution to (12) is computationally intractable due to the involved nonconvex optimization. To bypass this computational challenge, a local Gaussian approximation around $\bar{x}_{t-1}^{j}$ is employed such that
@@
This process is repeated until $t=\tau$ for each particle to obtain the set of trajectories $\{\bar{x}_{0:\tau}\}_{j=1}^{J}$, which will include the coarse estimates of local GNEs. This completes the unscented implicit particle filter, identifying the coarse estimates of multiple local Nash equilibria in (1).
In the last step, the local equilibria are extracted to be from $\{x_{0:\tau}\}_{j=1}^{J}$. As viewed in Fig. 4, the particle set contains multiple trajectories for each GNE. Therefore, to isolate the GNE trajectories, a clustering method is used to aggregate the particles close to each other in one cluster. One can leverage methods such as hierarchical clustering [35] or DBSCAN [36]. It should be noted that clustering methods require a metric to compute distances between the data points. The Fréchet distance [37] is utilized as our metric to compute distances between the trajectories. The clustering of trajectories will result in one cluster for each GNE by taking the average of trajectories over each cluster and feed them as a warm start to the IPOPT solver to obtain the local GNEs. The algorithmic steps of the proposed MultiNash-PF are provided in Algorithm 1.
@@
4:Initialize the particles: $\bar{x}_{0}^{j}$, $\hat{\Sigma}^{x,j}_{0}$, and $w_{0}^{j}$ for $j=1,\ldots,J$
5:Employ Algorithm-1 from [31] combined with UKF to obtain coarse solutions, $\{x_{0:\tau}\}_{j=1}^{J}$.
6:Extract trajectories through clustering from $\{x_{0:\tau}\}_{j=1}^{J}$ and warm start IPOPT solver to obtain the local GNEs $\{{x}_{0:\tau}^{i^{\star}},{u}_{0:\tau}^{i^{\star}}\}_{i=1}^{N}$ of $\mathcal{G}$.
@@ V Numerical Simulations
$p^{i}_{t+1}$ $=p^{i}_{t}+\Delta t\cdot v^{i}_{t}\cos{\theta^{i}_{t}},\;q^{i}_{t+1}=q^{i}_{t}+\Delta t\cdot v^{i}_{t}\sin{\theta^{i}_{t}}$
$\theta^{i}_{t+1}$ $=\theta^{i}_{t}+\Delta t\cdot\omega_{i},\;v^{i}_{t+1}=v^{i}_{t}+\Delta v^{i}_{t},\;\omega^{i}_{t+1}=\omega^{i}_{t}+\Delta\omega^{i}_{t},$
where $\Delta t$ is the time-step size, $p^{i}_{t}$ and $q^{i}_{t}$ are $x$ and $y$ coordinates of the positions, $\theta^{i}_{t}$ is the heading angle from positive x-axis, $v^{i}_{t}$ is the linear velocity, and $\omega^{i}_{t}$ is the angular velocity of the agent $i$ at time $t$. Further, $\Delta v^{i}_{t}$ is the change in linear velocity, and $\Delta\omega^{i}_{t}$ is the change in angular velocity of the agent $i$ at time $t$, respectively. For each agent $i$, we denote the state vector $x^{i}_{t}$ and the action vector $u^{i}_{t}$ as follows
$x^{i}_{t}=\begin{bmatrix}p^{i}_{t}&q^{i}_{t}&\theta^{i}_{t}&v^{i}_{t}&\omega^{i}_{t}\end{bmatrix}^{\top},\enskip u^{i}_{t}=\begin{bmatrix}\Delta v^{i}_{t}&\Delta\omega^{i}_{t}\end{bmatrix}^{\top}.$
We formulate this scenario as a two-player game according to cost functions as described in (3). The reference trajectory for both agents is given as a straight line on the highway, starting from their initial position to their respective goal location. With $\hat{Q}=diag(50,10,5,5,2)$, the reference tracking and control effort cost matrices are set as
@@
where $d(x^{1}_{t},x^{2}_{t})\coloneq\sqrt{(p^{1}_{t}-p^{2}_{t})^{2}+(q^{1}_{t}-q^{2}_{t})^{2}}$ is the Euclidean distance between the two agents and $r_{\text{col}}\coloneq 3m$ is the collision avoidance radius between the two agents.
@@
With this setup, we employ MultiNash-PF from Algorithm 1 to efficiently capture the multi-modality in the solution space and then extract the solutions using hierarchical clustering [35] to provide a warm start trajectory to IPOPT [24] solver to recover different local GNEs. We use the Julia programming language and utilize the JuMP [38] library to solve the game using the IPOPT solver. The results of our method are provided in Fig 2. As we can see, the presence of two agents and collision avoidance constraints induces two different possible GNEs. Either both the agents will yield to their right, or they will yield to their left to avoid collision. As shown in Fig 2, our method can recover both equilibria.
@@
Furthermore, we consider a more complicated scenario in which two agents aim to exchange their positions while avoiding an obstacle of radius $r_{\text{obs}}=4m$ located at the midpoint of the two agents. It should be noted that the radius of the obstacle is larger than the collision avoidance radius of the two agents.
@@
We consider the same unicycle dynamics as before, and the reference trajectory cost is also the same as in (13). Furthermore, the collision avoidance radius between the two agents is also the same as before. In addition to (14), the constraints for this game will include both agents avoiding the obstacle $g^{i}_{o}(x_{t},u_{t}):=\begin{bmatrix}-d(x^{i}_{t},x_{\text{obs}})+r_{\text{obs}}\end{bmatrix}$ for $i=1,2$ where $x_{\text{obs}}$ is the location of the obstacle. We also assume that we have constraints on the control bounds $g_{b}(x_{t},u_{t})=\begin{bmatrix}-v^{1}_{t},&-v^{2}_{t},&|u^{1}_{t}|-u^{1}_{b},&|u^{2}_{t}|-u^{2}_{b}\end{bmatrix}$, where $u^{1}_{b}$ and $u^{2}_{b}$ are controlled action bounds for agents 1 and 2, respectively. We write the combined constraints as $g(x_{t},u_{t}):=\begin{bmatrix}g_{c}&g^{1}_{o}&g^{2}_{o}&g_{b}\end{bmatrix}^{\top}\leq 0.$ We choose the values of control action bounds to be $u^{1}_{b}=u^{2}_{b}=[0.15,0.75]^{\top}$. This setup induces multitudes of local equilibria. Two trivial local equilibria are when the two agents pass each other through opposite sides of the obstacle. However, there exist four other nontrivial equilibria when both agents can go on the same side of the obstacle, and one of the agents can yield to the other one. It is challenging to recover all six equilibria without utilizing the capabilities of the implicit particle filter. Upon performing simulation experiments with our method, we observe that MultiNash-PF can recover all six Nash equilibria as shown in Fig. 4 and Fig. 3.
(a) Histogram of the solve time comparison of MultiNash-PF and the baseline.
(b) Histogram of the number of the IPOPT runs required to obtain all six equilibria.
@@
(a) The average solve time for MultiNash-PF is $3.36\pm 0.59s$ while the baseline’s average solve time is $6.71\pm 2.83s$. MultiNash-PF reduces the solve time by about $50\%$ compared to the baseline while giving a much lower variance on the solve time. (b) For the baseline, the average number of IPOPT runs required to obtain all six different equilibria is $18.64\pm 7.57$, while for MultiNash-PF, this is always $6$. MultiNash-PF requires 3 times fewer IPOPT runs.

### v3

@@ IV Identifying Local Nash Equilibria via Bayesian Inference
$\min_{\{v_{t},w_{t},\eta_{t}\}_{t=0}^{\tau}}$ $\left\lVert v_{\tau}\right\rVert_{Q_{\tau}}^{2}+\sum_{t=0}^{\tau-1}\left\lVert v_{t}\right\rVert_{Q}^{2}+\sum_{t=0}^{\tau}\left(\left\lVert w_{t}\right\rVert_{R}^{2}+\left\lVert\eta_{t}\right\rVert^{2}_{Q_{\eta}}\right),$
@@
where $v_{t}$, $w_{t}$, and $\eta_{t}$ are bounded disturbances. The weight matrix $Q_{\eta}\in\mathbb{R}^{n_{c}\times n_{c}}$ quantifies the slackness in satisfying constraints. Further, the degree of constraint violation is captured by a barrier function $\psi(\cdot,\cdot)$, which is defined as
@@
where $\alpha$ is a parameter that tunes the strictness of constraint enforcement.
A direct attempt at solving (IV) using traditional gradient-based methods poses a similar computational burden in identifying all of the local equilibria compared to problem (1). However, the above connection between (1) and the moving horizon estimation problem (IV) highlights the possibility to view the problem from a Bayesian inference perspective. To this end, we construct the following virtual state-space model from (IV) as
@@
with $\bar{Q}\coloneqq\mathop{\rm blkdiag}(Q^{-1},Q_{\eta}^{-1})$, $\bar{Q}_{\tau}\coloneqq\mathop{\rm blkdiag}(Q_{\tau}^{-1},Q_{\eta}^{-1})$, $\bar{R}\coloneqq\mathop{\rm blkdiag}(0_{n\times n},R^{-1})$. The state estimation relevant to (9) is to estimate the state $\bar{x}_{t}$ given the measurements $\bar{y}_{t}$. The value of the virtual measurements must be set in a way that steers the inference problem to have the same optima as in problem (1). We achieve this by noting that $y_{x,t}$ is a measurement value used to correct the state estimate of $x_{t}$ towards the reference $\hat{x}_{t}$ while $y_{g,t}$ is to satisfy constraints through (8). Observing that the barrier function in (8) outputs zero when the constraints are satisfied, we set the virtual measurement for the barrier function to be equal to $0_{c}$. Hence, the virtual measurement that will drive $\bar{x}_{t}$ towards optimality takes the value
@@
The Bayesian inference problem pertaining to (9) is one that characterize the conditional distribution $p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})$. In the following lemma, we show that an equivalence holds between a posteriori estimate of $p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})$ and (1).
@@
We refer the interested reader to [14, Thm. 1] for the proof. The above lemma implies that the local GNEs in (1) translate to multiple local maxima of $p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})$. Hence, the Bayesian inference problem for (9) involves characterizing the multi-modal PDF $p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})$. Generally, state estimation of a multi-modal PDF does not admit closed-form solutions [34]. Hence, we resort to approximate solutions. A powerful approach is to use Monte Carlo methods that empirically approximate the distribution via a set of $J$ particles as
@@
where $\delta(\cdot)$ is the Dirac delta function and $w_{\tau}^{i}$ is the weight assigned to a sample trajectory $\bar{x}_{0:\tau}^{i}$ for $j=1,\ldots,J$. It is important to note that, in our setting, each particle represents a complete trajectory.
Figure 3: MultiNash-PF identifies all the local GNE trajectories for a challenging two-agent trajectory game where two agents swapped positions while avoiding an obstacle. In this scenario, the obstacle radius is greater than the collision avoidance radius of the two agents, leading to the discovery of 6 distinct Nash equilibria. MultiNash-PF effectively obtains coarse estimates of local GNEs through implicit particle filtering, which are then refined using IPOPT to obtain local GNEs.
To achieve the approximation in (10), we employ the implicit particle filter in [31] due to the twofold benefit it holds. First, as suggested in Lemma 1, the local maxima (i.e., high-probability regions) of $p(\bar{x}_{0:\tau}\,|\,\bar{y}_{0:\tau})$ correspond to the local GNEs. Hence, focusing our sampling efforts towards these high-probability regions can reduce the number of required samples for the approximation in (10). This approach aligns with the principle of implicit importance sampling, whose objective is to build a mapping that draws particles from the high-probability regions of the filtering distribution in (10). The implicit particle filter mitigates the particle degeneracy problem that pertains to particle filters by focusing the samples on the high-probability regions, which results in samples with large enough weights that are less prone to degeneration. Therefore, the implicit particle filter can often be deployed without requiring a resampling step [31], thereby resulting in smoother approximated trajectories, although occasional resampling may still be required.
We leverage the Markovian property of the virtual system (9) to recursively approximate (10) using Bayes’ rule:
@@
for $0\leq t\leq\tau$. This indicates that given the particle $\bar{x}_{t-1}^{j}$ from the prior distribution $p(\bar{x}_{0:t-1}\,|\,\bar{y}_{0:t-1})$, we can obtain $\bar{x}_{t}^{j}$. As shown in [31], we use implicit importance sampling and identify an implicit mapping to sample from high probability regions of (11). The implicit map is constructed by considering sampling from a reference random vector $\xi_{t}$ with density $p(\gamma_{t})$ and mapping it to high probability regions of $p(x_{0:\tau}\,|\,y_{0:\tau})$, such that a sample $\gamma^{i}_{t}$ is mapped to $x_{t}^{i}$. The mapping connects the highly probable regions of $p(\xi)$ to $p(x_{0:\tau}\,|\,y_{0:\tau})$ by letting
@@
where $F(x_{t}^{i})=-\log\left(p(x_{t}^{i}\,|\,y_{1:t})\right)$ and $G(\gamma_{t}^{i})=-\log\left(p(\gamma_{t}^{i})\right)$. Finding an explicit solution to (12) is computationally intractable due to the involved nonconvex optimization. To bypass this computational challenge, we employ a local Gaussian approximation around $\bar{x}_{t-1}^{j}$ such that
@@
We repeat this process until $t=\tau$ for each particle to obtain the set of trajectories $\{\bar{x}_{0:\tau}\}_{j=1}^{J}$, which will include the coarse estimates of local GNEs. This completes the unscented implicit particle filter, which identifies the coarse estimates of multiple local Nash equilibria in (1).
In the last step, we extract local equilibria from $\{x_{0:\tau}\}_{j=1}^{J}$. As viewed in Fig. 2, the particle set contains multiple trajectories for each GNE. Therefore, to isolate the GNE trajectories, we use a clustering method to aggregate the particles close to each other in one cluster. We leverage methods such as hierarchical clustering [36] or DBSCAN [37] for clustering purposes. It should be noted that clustering methods require a metric to compute distances between the data points. We use Fréchet distance [38] as our metric to compute distances between the trajectories. The clustering of trajectories will result in one cluster for each GNE. For each cluster, we compute the average trajectory which serves as an initial guess for the IPOPT solver. The solver then refines this initial guess by minimizing (1), ultimately converging to a precise local GNE. The algorithmic steps of the proposed MultiNash-PF are provided in Algorithm 1.
@@
[…]
6:  Employ Algorithm-1 from [31] combined with UKF
7:  to obtain coarse solutions, $\{x_{0:\tau}\}_{j=1}^{J}$.
8:Extract trajectories through clustering from $\{x_{0:\tau}\}_{j=1}^{J}$ and warm start IPOPT solver to obtain the local GNEs $\{{x}_{0:\tau}^{i^{\star}},{u}_{0:\tau}^{i^{\star}}\}_{i=1}^{N}$ of $\mathcal{G}$.
@@ V Numerical Simulations
$p^{i}_{t+1}$ $=p^{i}_{t}+\Delta t\cdot\nu^{i}_{t}\cos{\theta^{i}_{t}},\;q^{i}_{t+1}=q^{i}_{t}+\Delta t\cdot\nu^{i}_{t}\sin{\theta^{i}_{t}}$
$\theta^{i}_{t+1}$ $=\theta^{i}_{t}+\Delta t\cdot\omega_{i},\;\nu^{i}_{t+1}=\nu^{i}_{t}+\Delta\nu^{i}_{t},\;\omega^{i}_{t+1}=\omega^{i}_{t}+\Delta\omega^{i}_{t},$
where $\Delta t$ is the time-step size, $p^{i}_{t}$ and $q^{i}_{t}$ are $x$ and $y$ coordinates of the positions, $\theta^{i}_{t}$ is the heading angle from positive x-axis, $\nu^{i}_{t}$ is the linear velocity, and $\omega^{i}_{t}$ is the angular velocity of the agent $i$ at time $t$. Further, $\Delta\nu^{i}_{t}$ is the change in linear velocity, and $\Delta\omega^{i}_{t}$ is the change in angular velocity of the agent $i$ at time $t$, respectively. For each agent $i$, we denote the state vector $x^{i}_{t}$ and the action vector $u^{i}_{t}$ as follows
$x^{i}_{t}=\begin{bmatrix}p^{i}_{t}&q^{i}_{t}&\theta^{i}_{t}&\nu^{i}_{t}&\omega^{i}_{t}\end{bmatrix}^{\top},\enskip u^{i}_{t}=\begin{bmatrix}\Delta\nu^{i}_{t}&\Delta\omega^{i}_{t}\end{bmatrix}^{\top}$.
We formulate this scenario as a two-player game according to cost functions as described in (3). The reference trajectory for both agents is given as a straight line on the highway, starting from their initial position to their respective goal location. With $\hat{Q}=\mathop{\rm blkdiag}(50,10,5,5,2)$, the reference tracking and control effort cost matrices are set as
@@
where $d(x^{1}_{t},x^{2}_{t})$ is the Euclidean distance between the two agents and $r_{\text{col}}\coloneq 3m$ is the collision avoidance radius between the two agents.
@@
With this setup, we employ MultiNash-PF from Algorithm 1 to efficiently capture the multi-modality in the solution space and then extract the solutions using hierarchical clustering [36] to provide a warm start trajectory to IPOPT [24] solver to recover different local GNEs. We use the Julia programming language and utilize the JuMP [39] library to solve the game using the IPOPT solver. The results of our method are provided in Fig 4. As we can see, the presence of two agents and collision avoidance constraints induces two different possible GNEs. Either both agents will yield to their right, or they will yield to their left to avoid collision. As shown in Fig 4, our method can recover both equilibria.
Figure 4: The Nash equilibrium trajectories found by MultiNash-PF when two unicycle agents change their positions. First, we use the implicit particle filter to discover the coarse estimates of two different equilibria, and then utilize them as a warm start for IPOPT to obtain the two Nash equilibrium trajectories.
@@
Next, we consider a more complicated scenario in which two agents aim to exchange their positions while avoiding an obstacle of radius $r_{\text{obs}}=4m$ located at the midpoint of the two agents. Note that the radius of the obstacle is larger than the collision avoidance radius of the two agents.
@@
We consider the same setup as before. In this example, in addition to (14), the constraints for this game will include both agents avoiding the obstacle $g^{i}_{o}(x_{t},u_{t}):=\begin{bmatrix}-d(x^{i}_{t},x_{\text{obs}})+r_{\text{obs}}\end{bmatrix}$ for $i=1,2$ where $x_{\text{obs}}$ is the location of the obstacle. We also assume that we have constraints on the control bounds $g_{b}(x_{t},u_{t})=\begin{bmatrix}-\nu^{1}_{t},&-\nu^{2}_{t},&|u^{1}_{t}|-u^{1}_{b},&|u^{2}_{t}|-u^{2}_{b}\end{bmatrix}$, where $u^{1}_{b}$ and $u^{2}_{b}$ are controlled action bounds for agents 1 and 2, respectively. We write the combined constraints as $g(x_{t},u_{t}):=\begin{bmatrix}g_{c}&g^{1}_{o}&g^{2}_{o}&g_{b}\end{bmatrix}^{\top}\leq 0.$ We choose the values of control input bounds to be $u^{1}_{b}=u^{2}_{b}=[0.15,0.75]^{\top}$. This setup induces multitudes of local equilibria. Two trivial local equilibria are when the two agents pass each other through opposite sides of the obstacle. However, there exist four other nontrivial equilibria when both agents can go on the same side of the obstacle, and one of the agents can yield to the other one. It is challenging to recover all six equilibria without utilizing the capabilities of the implicit particle filter. We fixed $J=50$ particles and obtained the solution trajectories using implicit particle filtering from Algorithm 1 described in Section IV. Upon performing simulation experiments with our method, we observe that MultiNash-PF can recover all six Nash equilibria as shown in Fig. 2 and Fig. 3. Please note that, while trajectories may overlap spatially, collisions are avoided because agents traverse those regions at different times, as enforced by the constraints in our optimization.
(a) Histogram of the number of IPOPT runs required to obtain all six equilibria.
(b) Histogram of the solve time comparison of MultiNash-PF and the baseline.
@@
(a) For the baseline, the average number of IPOPT runs required to obtain all six different equilibria is $18.64\pm 7.57$, while for MultiNash-PF, this is always $6$. MultiNash-PF requires 3 times fewer IPOPT runs. (b) MultiNash-PF reduces the solve time by about $50\%$ compared to the baseline while giving a much lower variance on the solve time.